OpenStudy (lgbasallote):

10*9*8*7*6*5*4*3*2*1 = 10! Can this be true?! Why or why not? hihi @FoolForMath >:)))

OpenStudy (diyadiya):

Yes ,Why not ?

OpenStudy (anonymous):

Of course its true

OpenStudy (lgbasallote):

lol =)) prove :ppp

OpenStudy (anonymous):

Depends on the definition of 10! and what do you mean by the operation of '*'

OpenStudy (lgbasallote):

those are multiplication symbols

OpenStudy (ash2326):

Then this is true:)

OpenStudy (mimi_x3):

lol, use a calculator.

OpenStudy (ash2326):

10*9*40320

OpenStudy (unklerhaukus):

both sides = 3628800

OpenStudy (anonymous):

Then this is true. 10! is 3628800 and the L.H.S 10*9*8*7*6*5*4*3*2*1 = 3628800 hence QED.

OpenStudy (anonymous):

Silly problem! NEXT :P

OpenStudy (lgbasallote):

oh...okay..i thought that was hard...

OpenStudy (mimi_x3):

lol, "hard" even a 5 year old can do it. :P

OpenStudy (anonymous):

Mimi is super amazing 10 year old kid ;)

OpenStudy (anonymous):

a*

OpenStudy (anonymous):

\[n!= \prod_{k=1}^{n}k=1*2*3*...*n\] This is the definition which would prove why it's correct.

OpenStudy (callisto):

@Mimi_x3 not really... at least my sis won't know, she hasn't learnt factorial yet!

OpenStudy (ash2326):

Guys I think it's becoming a medal party!!!!!

OpenStudy (anonymous):

\[n \in \mathbb{N}\]

OpenStudy (unklerhaukus):

it is good to see some Pi notation @Noliec

OpenStudy (mimi_x3):

@Callisto I'm sure that your sister is able to use a calculator.

OpenStudy (callisto):

Yes, even so, she won't know what is the right button :(